OpenStudy (kainui):

What is the general formula for the "tens place" of 3^n where n is greater than or equal to 0? It starts as 0,0,0,2,8,...

OpenStudy (compassionate):

Gonna have to think about this one,

OpenStudy (kainui):

Here's a hint: it's periodic. =P

Parth (parthkohli):

I'm thinking binomial theorem or modular arithmetic. Not like I've studied either of them...

OpenStudy (kainui):

Me either

OpenStudy (mathmath333):

\(\large \color{black}{ \normalsize \text{Ten's place can be found by}\hspace{.33em}\\~\\ 3^{n}\pmod {100} \hspace{.33em}\\~\\~\\~\\ \normalsize \text{or by binomial }\hspace{.33em}\\~\\ 3^{n}\hspace{.33em}\\~\\ \implies 9^{n/2}\hspace{.33em}\\~\\ \implies (-1+10)^{n/2}\hspace{.33em}\\~\\ \implies \dbinom{\small{\frac{n}{2}}}{0}(-1)^{n/2}(10)^{0}+\dbinom{\small{\frac{n}{2}}}{1}(-1)^{n/2-1}(10)^{1}\hspace{.33em}\\~\\ }\)

Parth (parthkohli):

@mathmath333 Should that not be \((3^n \% 100) - (3^n \% 10)\)?

Parth (parthkohli):

^ and what prevents that from being interpreted as the "general" formula? Huehuehue...

OpenStudy (mathmath333):

if ur finding \(3^{n}\pmod {100} \hspace{.33em}\\~\\\) then finding \(3^{n}\pmod {10} \hspace{.33em}\\~\\\) is just an extra work \(3^{n}\pmod {100} \hspace{.33em}\\~\\\) will give u ten's place as well as unit's place

OpenStudy (compassionate):

Upm/r1r2=565374.9999 We can all go home now

Parth (parthkohli):

but you want the tens place only. and for that, the formula would be (3^n % 100 - 3^n % 10)/10

OpenStudy (kainui):

I suppose you can consider \[\Large \frac{3^n \mod 100 - 3^n \mod 10 }{10}\] to be the general formula, but it's not really pretty imo.

OpenStudy (mathmath333):

if u want only ten's place u can find \(3^{n}\pmod {100} \hspace{.33em}\\~\\\) and then smartly ignore the last digit , no need to waste time on calculating \(3^{n}\pmod {10} \hspace{.33em}\\~\\\)

OpenStudy (kainui):

Haha that's sort of like a function that's the opposite of 1 to 1, it's a purposefully 10 to 0 function =P

Parth (parthkohli):

0 0 2 8 4 2 8 6 8 4 4 4 2 6 0 2 6 8 6

OpenStudy (kainui):

I think this is on the right path, now we know the period is 20.

Parth (parthkohli):

Another observation: these numbers correspond to the tens place of \(3^n\) :O Mindblown.jpg

OpenStudy (kainui):

I actually like the idea of ignoring the last digit... It's kind of a fascinating idea to just throw information away but maybe in a more formal sense. Hmm.

OpenStudy (kainui):

Wait, I thought that was the observation.

OpenStudy (kainui):

What were you thinking when you found these?

Parth (parthkohli):

How do we generate these? Another exponential function?

Parth (parthkohli):

I missed one zero in the beginning. Anyway...

OpenStudy (mathmath333):

@kainui are u trying to solve the ganeshie8's question on triangle perimeter

OpenStudy (kainui):

That's originally how I came across this problem, yes.

OpenStudy (kainui):

But I don't think it is really the best way to solve that problem, and it's its own problem to itself. =P

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} \normalsize \text{then this formula might help u}\hspace{.33em}\\~\\ \implies \dbinom{n/2}{0}(-1)^{n/2}(10)^{0}+\dbinom{n/2}{1}(-1)^{n/2-1}(10)^{1}\hspace{.33em}\\~\\ +\dbinom{n/2}{2}(-1)^{n/2-2}(10)^{2}+\dbinom{n/2}{3}(-1)^{n/2-3}(10)^{3}\hspace{.33em}\\~\\ \end{align}}\) as he asked about the remainder for \(3^n \pmod{10^4}\)

OpenStudy (perl):

*

OpenStudy (kainui):

Just to mix things up, if \[\Large 3^n = 3^{4a+b}, \ \ \ 0 \le b \le 3\] then\[\Large 10 \lfloor \frac{3^n}{10}\rfloor =8^{b+1}(a-1) \mod 10\] Sadly this is not true for when b=3. But maybe you'll find this interesting to keep you looking for the answer. ;P

OpenStudy (anonymous):

latex tip\[a\stackrel{c}{\equiv}b\]:-)